Machiel Manufacturing acquired a $60,000 machine on January 1, 2009. The machine is estimated to have a useful life of 4 years, and a residual value of $10,000. For units-of-production depreciation purposes, the machine is expected to produce 500,000 units. If Machiel Manufacturing uses double-declining-balance depreciation, what is the depreciation expense in 2010?

Select one: A. $25,000 B. $ 8,640 C. $ 9,600 D. $15,000 E. $12,500
Answer #1

Depreciation rate as per straight line method=100%/4=25%/year

Hence depreciation as per double decline balance=2*Depreciation rate as per straight line method*Beginning value of each period

Year Beginning value Depreciation Ending value
2009 60,000 (2*25%*60,000)=30,000 (60,000-30,000)=30,000
2010 30,000 (2*25%*30,000)=$15,000

Hence depreciation expense in 2010=$15,000